OpenStudy (anonymous):

integral of dy/sqrt(y)(1-sqrt(y))

OpenStudy (anonymous):

\[\int\limits \frac{ dy }{ \sqrt{y}(1-\sqrt{y}) }\]

OpenStudy (freckles):

let u=1-sqrt(y)

OpenStudy (anonymous):

how do i solve for du in terms of dy then?

OpenStudy (freckles):

are you asking how to differentiate sqrt(y)?

OpenStudy (anonymous):

no

OpenStudy (freckles):

you differentiate both sides

OpenStudy (freckles):

try

OpenStudy (anonymous):

i need to have du in the equation to solve the integral

OpenStudy (freckles):

you will end getting something like du=f(y) dy

OpenStudy (anonymous):

so then it equals \[\int\limits \frac{ dy }{ u+1(u) }\]

OpenStudy (kainui):

I picked sqrt(y)=u it seemed to work pretty well for me.

OpenStudy (freckles):

\[\int\limits\limits \frac{ dy }{ \sqrt{y}(1-\sqrt{y}) } \\ \int\limits_{}^{} \frac{-2}{1-\sqrt{y}} \cdot \frac{1}{-2 \sqrt{y}} dy\]

OpenStudy (anonymous):

\[\frac{ du }{ dy }=\frac{ 1 }{ 2\sqrt{y} }\]

OpenStudy (anonymous):

with a negtaive

OpenStudy (freckles):

if you choose u=1-sqrt(y)

OpenStudy (freckles):

\[u=1-\sqrt{y} \\ du=-\frac{1}{2 \sqrt{y}} dy\]

OpenStudy (kainui):

\[\LARGE \int\limits \frac{dy}{\sqrt{y}-y}\] \[\LARGE \sqrt{y}=u \\ \LARGE \frac{dy}{2 \sqrt{y}}=du \\ \LARGE dy=2udu\] \[\LARGE \int\limits \frac{2u du}{u-u^2}=-2 \int\limits \frac{du}{u-1}\\ \LARGE -2\ln|u-1|+C = \ln \frac{1}{(\sqrt{y}-1)^2}+C\]

OpenStudy (anonymous):

well then \[-2\ln \left| 1-\sqrt{y} \right|+C \] is my answer